OpenStudy (anonymous):

Let triangle A B C is similar to triangle D E F. In triangle a b c, side a b equals five and side b c equals ten. side d e is unknown, and side e f equals eight. What is the measure of side d e ? 4 8 10 16

OpenStudy (anonymous):

the side ab is congruent to side de side bc is congruent to side ef therefore: ab/bc=de/ef 5/10=x/8 now cross multiply can you do it now?

OpenStudy (anonymous):

ratio of their corresponding sides are in proportion so